OpenStudy (anonymous):

Find the percent discount.. regular price: $49 sale price: $39

hero (hero):

\[\% \Delta = \frac{|a-b|}{a} \times 100\] a = original price b = sale price

hero (hero):

\[\% \Delta = \frac{49- 39}{39} \times 100\]

hero (hero):

\[\% \Delta = \frac{10}{39} \times 100\]

hero (hero):

\[\% \Delta = \frac{1000}{39} = 25.641\approx 26\]

hero (hero):

approximately 26% change